OpenStudy (anonymous):

Find the diagonal of a cube if its side equals 5. help

OpenStudy (anonymous):

diagonal = sqrt 3 times a

OpenStudy (anonymous):

 a = measure of the cube side

OpenStudy (anonymous):

Let s = length of the side of a cube Therefore, using the Pythagorean theorem, s^2 + s^2 = 6^2 2s^2 = 36 s = 6/sqrt 2 Let D = diagonal of the cube Again, using the Pythagorean theorem, D^2 = 6^2 + (6/sqrt 2)^2 D^2 = 36 + 36/2 = 36 + 18 = 54 D = sqrt (54) = sqrt (9 * 6) D = 3*sqrt (6) units Hope this helps.

OpenStudy (anonymous):

diagonal of cube is \[\sqrt{3}\] times side therefore diagonal = 5 \[\sqrt{3}\]

OpenStudy (anonymous):